the woods, walking 1.46 km due east, 0.519 km due south, and 1.60 km in a direction 46.0 north of west. The time required for this trip is o.8687 h. Determine the magnitudes of the bird watchers (a) displacement and (b) average velocity. 19.569 1.7836 MacBook Air
Answer #1

a] displacement = (1.46 - 1.6 cos 46 degree ) i + ( 1.6 sin 46 degree - 0.519)j

= 0.34854i + 0.6319 j

= sqrt(0.6319^2 +0.34854^2)

= 0.7216 km answer

b] v = d/t = 0.7216/0.8687 = 0.8307 km/h answer